OpenStudy (anonymous):

Prove that sqrt((y^2/a^2)+1) is a catenary

OpenStudy (anonymous):

\[\sqrt{(y^2/a^2)+1}=x\]

OpenStudy (anonymous):

i was thinking of doing dx/dy

OpenStudy (anonymous):

What do you have about the mathematical definition of a catenary?

OpenStudy (anonymous):

I'd say that's a step in the right direction.

OpenStudy (anonymous):

\[x^2=(y^2/a^2)+1\] \[y^2=a^2(x^2+1)\] catenary is cosh(u)

OpenStudy (anonymous):

\[2y * y \prime =2a^2x\] \[y \prime = 2a^2x / 2y\] \[y \prime = 2a^2x/ \sqrt{a^2(x^2+1)}\] i'm not too sure if i can take the a^2 out of the sqrt to cancel it.

OpenStudy (anonymous):

I think you'll want to use this definition \[y = \frac{e^{ax} +e^{-ax}}{2a}\] Where a is a constant.

OpenStudy (anonymous):

@SmoothMath , ill check that out. thanks!

OpenStudy (anonymous):

This explains it well, and I'm feeling pretty sure that's how to do it. I'll see if I can formalize it. http://www.proofwiki.org/wiki/Catenary

OpenStudy (anonymous):

\(x = \sqrt{(y^2/a^2)+1} \) \(x^2 =y^2/a^2 +1 \) \(x^2 - 1 = y^2/a^2 \) \(y^2 = (x^2-1)*a^2\) \( y = \sqrt{(x^2-1)*a^2} = \sqrt{x^2-1}\sqrt{a^2} = \sqrt{x^2-1}*a\)

OpenStudy (anonymous):

hmph...

OpenStudy (anonymous):

or maybe this is the route to go... \(\sqrt{x^2a^2-a^2}\)

OpenStudy (anonymous):

\[y \prime = 2a/\sqrt{x^2+1} \] \[\int\limits_{}^{} y \prime = 2a \int\limits_{}^{}1/ \sqrt{x^2+1} \] \[y = 2a*sinh^{-1} x \] wtf. now it's in sine. and sine is not a catenary. hahahhaa i'll try to continue.